OpenStudy (anonymous):

find all roots x^4-3x^2-10=0

OpenStudy (asnaseer):

substitute \(u=x^2\) and solve the quadratic in u

OpenStudy (asnaseer):

then solve for x

OpenStudy (anonymous):

Factoring -10=-5(2) (X^2-5)(x^2+3)=0 x^2-5=0 or x^2+3=0 And you solve for x